OpenStudy (anonymous):

Write an EQUATION for the line that passes THROUGH the points (0, -6) and (-3, 0). a:y= -2x- 6 b:y= -2x- 3 c:y= -1/2x - 6 d:y= -1/2x - 3

OpenStudy (anonymous):

@KeganMoore @Nurali @waterineyes @ganeshie8 help?

OpenStudy (anonymous):

@amistre64 @Hero @jim_thompson5910

OpenStudy (anonymous):

Use two point slope form..

OpenStudy (anonymous):

\[\frac{y - y_1}{y_2 - y_1} = \frac{x - x_1}{x_2 - x_1}\]

OpenStudy (anonymous):

Just plug in the values there..

OpenStudy (anonymous):

please someone explain me whats slope i'm completely lost!

OpenStudy (anonymous):

\[x_1 = 0, x_2 = -3\] \[y_1 = -6, y_2 = 0\]

OpenStudy (anonymous):

\[Slope = \frac{Difference \; \; of \; \; y \;\;coordinates}{Difference \;\;of \;\; x \;\;coordinates}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

...

OpenStudy (anonymous):

Here, can you find slope??

OpenStudy (anonymous):

Firstly tell me what is the difference of y coordinates there?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

-6?

OpenStudy (anonymous):

So, after finding slope write here, then we will guide you further..

OpenStudy (anonymous):

See, you have done -6 - 0 to get -6. In the same way you will find difference of x coordinates.

OpenStudy (anonymous):

x = -3 y = -6 ?

OpenStudy (anonymous):

the difference...

OpenStudy (anonymous):

What is difference of x one??

OpenStudy (anonymous):

I mean what you got for difference of x coordinate?

OpenStudy (anonymous):

it's -3?

OpenStudy (anonymous):

0 - (-3) = ??

OpenStudy (anonymous):

so it's "c"

OpenStudy (anonymous):

Why it can't be d??

OpenStudy (anonymous):

"c" is pretty good..

OpenStudy (anonymous):

u think it's d or c u know this better then I do

OpenStudy (anonymous):

Ha ha ha.. I think none.. You said "c", so I just asked you why it can't be d? (may be you will be guessing also).. Then, I check which one it is, then I found c..

OpenStudy (anonymous):

Answer part matters to you not me, I want to give just the method of finding it.. :)

OpenStudy (anonymous):

ok! Thanks

OpenStudy (anonymous):

You are welcome dear..